Q- If a ideal wire (having no resistance) is connected to a DC power supply of Voltage(V), what will be the current in the wire?


And if current i=0, is the potential difference still V?


 

Grade:9

2 Answers

Vaishal Shah
13 years ago

Dear Shivam,

If R=0 then,

according to Ohm's law,

current(I)=V/R

             =V/0

             =Infinity

 

And if i=o then,

V=I*R

 =0.

iit jee
13 years ago

acc. to ohms law v=ri

if r=o then current will be infinity.

v=ri.

if i=0 then v=zero........

plzzz approve!!!!!!!!

 

kedar joshi.